Solve by completeing the square.
x squared - 6x -3 = 0


2x squared + 10x + 11 = 0

Answer by checkley71(8403)

X^2-6X-3=0
X^2-6X+9=3+9
(X-3)^2=12
X-3=SQRT12
X=3+3.464
X=6.464 ANSWER.
----------------------------------
2X^2+10X+11=0
2X^2+10X=11
2(X^2+5X+6.26)=11+2*6.25
2(X+5/2)^2=11+12.5
(X+5/2)^2=23.5/2
(X+5/2)^2=11.75
X+2.5=SQRT11.75
X=-2.5+3.4278
X=.92378 ANSWER.